33. Consider the following reaction : CO(g)+2H 2 (g) CH 3 OH(liq.) Given : Delta f G^ (CO)=-137 kJ.mol^ -1 , Delta f G^ (CH 3 OH)=-166 kJ.mol^ -1 Calculate Delta_{r}*G at 25 degrees * C for the following pressures : P(CO) = 5 atm and P(H_{2}) = 3atm .

1
Expert's answer
2022-04-06T11:35:04-0400

∆G0rxn = "\Sigma"∆Gprod - "\Sigma"∆Greact = ∆G(CH3OH) - (∆G(CO) + 2×∆G(H2)) = -166 kJ/mol - (-137 kJ/mol + 2×0 kJ/mol) = -29 kJ

∆Grxn = ∆G0rxn + RTlnQ

Q = 1 / (pCO × p2H2) = 1 / (5 atm × (3 atm)2) = 0.022

∆Grxn = -29 kJ + (0.00831 kJ/mol K) × (298 K) × ln(0.022) = -38.45 kJ